OpenStudy (anonymous):

sin^2(3x)=1−cos6x , can someone explain the process?

OpenStudy (anonymous):

sin^2(3x) = 1 - cos^2(3x) might be useful here

OpenStudy (anonymous):

no start with cos 6x = 1 - 2sin^2(3x) 1 - cos 6x = 2sin^2(3x) - are you sure you have the correct expression

OpenStudy (anonymous):

Jimmyrep had the right idea...

OpenStudy (anonymous):

\[1-\cos(6x)=2\sin^23x\]

OpenStudy (anonymous):

Oh .

OpenStudy (anonymous):

Not so bright is correct...to sub

OpenStudy (anonymous):

\[\sin^23x=0\] \[3x=n \pi\]

OpenStudy (anonymous):

yes i think notsobright is correct

OpenStudy (anonymous):

cos(2u)=1-2sin^2(u)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

since u = 3x then cos2(3x)= 1-2sin^2(3x)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

this is what jimmyrep had to start with...not what is given...make sure you copied it down correctly.

OpenStudy (anonymous):

Alright .

OpenStudy (anonymous):

notsobright is correct

OpenStudy (anonymous):

I first thought he/she were trying to verify that they were equal.... but as notsobright and kingkos has indicated...it appears that you were solving for x in which case not so bright was correct...

OpenStudy (anonymous):

values in which the sin^2(3x)=0...or when 3x=npi as was stated. sorry for the confusion.

OpenStudy (anonymous):

Got it , thanks again !